Mr khan won a cash prize of certain amount int he jackpot of a carnival. He gave half of his money and an additional money of Rs 10000 to his wife. He also gave one- third of his money and an additional money of Rs 3000 to his son. If his wife got double the amount his son got find the amount of money Mr khan won in the jackpot
Answer by addingup(3677)

Certain amount: x
1/2x+10000+1/3x+3000 = x
1/2x+1/3x=13000 = x
3/6x+2/6x = x-13000
5/6x = x-13000
5/6x-x = -13000
5/6x-6/6x = 13000
-1/6x = --13000
x = -13000*-6 = 78000 (remember -x*-x = +x)
``````````````````````````````````````````````````